Each of two groups consists of 100 patients who have leukaemia. A new drug is given to the first group but not to the second (the control group). It is found that in the first group 60 people have remission for 2 years; but only 40 in the second group. Find 95% CI.
Answer by Boreal(15235)

two sample proportion 95% half-interval is z*SE
SE: start with group 1 SE which is sqrt(0.6*0.4/100) or 0.0490, and group 2 SE which is the same
the SE is sqrt(0.0490^2+0.0490^2)=0.0693
z*SE=1.96*0.0693=0.1358
the actual 95%CI is the difference +/- SE which is (0.0642, 0.3358). We are looking at the difference in proportion of survival between 1 and 2.
This is consistent with the highly significant difference.